### v2

# Reduced-Order Hydrodynamic Modelling of a Sphere Near a Wall Using Sparse Regression and Neural Operators

Abstract
This work presents an interpretable parametric surrogate model motivated by the need to identify a hydrodynamic model for resolving the trajectory of an object in real-time. The surrogate is formulated as a reduced-order model for a canonical configuration in which a one-degree-of-freedom heaving sphere operates near a vertical wall. High-fidelity CFD simulations are used to generate a parametric dataset of heave-decay responses over varying wall distances (WD) and drop heights (DH). Sparse Identification of Nonlinear Dynamics (SINDy) is then applied to each CFD trajectory to identify a low-order nonlinear ordinary differential equation (ODE) with polynomial terms representing effective hydrostatic restoring and radiation damping, and the harmonic terms representing the wave-induced excitation forces. The SINDy-identified coefficients are then used as a prior constraint in a neural operator network (ONet) that learns a smooth mapping from wall distance and drop height to the ODE coefficients, yielding a surrogate capable of predicting dynamics at arbitrary points in the input space without rerunning expensive CFD calculations. The resulting surrogate reproduces CFD heave-decay responses with near-optimal accuracy given the limiting assumptions while being capable of running in real time. The approach provides a practical pathway toward real-time, physics-informed surrogate modelling for launch-and-recovery operations.
Introduction
Launch and recovery of small uncrewed surface vessels (USVs) from larger parent vessels is a key enabler for future naval operations, but it remains one of the most challenging phases of the mission. During launch and recovery, the USV is manoeuvred in close proximity to a much larger vessel, often in steep and irregular seas.
Conventional tools for predicting ship motions, such as strip theory and linear potential-flow methods, are well established and computationally efficient under the assumptions of inviscid, irrotational flow and small-amplitude motions [19]. In this linear regime, hydrodynamic forces are commonly decomposed into linear added mass, radiation damping, hydrostatic stiffness, and first-order wave-excitation components [28, 9]. However, as wave steepness, motion amplitude, and geometric complexity increase, this linear description becomes progressively less accurate: higher-order wave interactions, non-linear wave–radiation, and non-linear hydrostatic effects are no longer adequately captured [24, 18].Similar limitations have been reported for ship–ship interactions in close proximity, where wave resonance in gaps and strong reflection effects introduce additional non-linearities [26, 32].
To resolve these effects with higher fidelity, recent studies have turned to Computational Fluid Dynamics (CFD), which can represent viscous effects, and non-linear free-surface dynamics at the expense of substantially increased computational cost [14].
However, because high-fidelity CFD remains too costly for real-time use, recent work in naval hydrodynamics has increasingly turned to surrogate modelling strategies capable of reproducing key nonlinear behaviours at a fraction of the computational expense. Data-driven sequence models including LSTM [10, 17, 27, 29, 25], and hybrid CNN–RNN architectures [5, 16, 30, 22] have been increasingly applied to predict ship motions in waves showing some promising results for accurate real-time short-term motion prediction [6, 23]. Modal-decomposition approaches such as POD and DMD have also been used to construct low-dimensional representations that can be combined with neural networks for nonlinear forecasting [1, 7]. More recently, physics-informed (PINNs) [20, 8] and operator-learning methods (Onets) [31], have been introduced to wave–body interaction problems to improve generalisation in data-sparse regimes. In parallel, sparse-regression techniques such as SINDy have shown that compact and interpretable nonlinear ODEs can be recovered directly from time series dynamics data [3, 12] .
Despite this progress, most existing surrogate models focus on isolated single-body dynamics. Comparatively little work has addressed the multi-scale, multi-body interactions characteristic of launch-and-recovery operations [4, 32], motivating the need for physics-interpretable surrogates such as the one developed in this study.
Background
Simplified LARS Configuration
In previous work, the present authors developed and validated a CFD framework for a simplified LARS scenario in which a small USV is idealised as a one-degree-of-freedom (1-DoF) heaving sphere operating near a large vertical wall representing the mothership shown in Figure 1 and 2. This idealised problem is chosen because it retains the key hydrodynamic mechanisms relevant to launch and recovery while remaining well-supported by prior studies, including experimental measurements of sphere heave-decay motion [14], responses of a sphere under regular wave forcing [18], and analytically derived solutions [13].
Figure 1: Schematic of the heaving sphere near a vertical wall, representing a simplified LARS configuration.
Figure 2: Freeze frame from CFD simulation.
CFD-Based Parametric Dataset
Previous work by this team [11] have demonstrated that CFD can capture key nonlinear hydrodynamic through comparing the simulation output with benchmark experimental data and analytical solutions. Based on the validated flow model, a parametric dataset of heave-decay responses across wall distances $WD\in(9,25)$m and drop heights $DH\in(0.5,5)$m was constructed. The drop heights considered are intentionally kept less than the sphere’s radius to ignore physics relating to an entry-type problem. The heave response dataset, depicted in Figure 3, is normalised by $\frac{Hv}{WD}$. A strong sensitivity on the input conditions (WD,DH) is evident.
Figure 3: Raw dataset; train-test split
Motivation
While CFD is an appropriate tool for resolving the underlying physics of USV launch and recovery, it remains too expensive to employ directly for real-time usage. A single 30 s heave decay simulation dropping a sphere into water can require $\mathcal{O}(10^{2}-10^{3})$ CPU core-hours, and a comprehensive parametric study over WD and DH would demand thousands of such runs.
Methodology
High Fidelity Hydrodynamic Modelling
Ansys Fluent 2025R1 is used to numerically solve the Navier-Stokes (Equations 1, 2) for an isothermal and incompressible flow on a structured Cartesian grid.
$\nabla\cdot\boldsymbol{u}$ $=0,$ (1)
$\frac{\partial\boldsymbol{u}}{\partial t}+\left(\boldsymbol{u}\cdot\nabla\right)\boldsymbol{u}$ $=-\frac{1}{\rho}\nabla p+\nu\nabla^{2}\boldsymbol{u}+\boldsymbol{g},$ (2)
A dynamic overset mesh with 1-DoF heave motion is updated explicitly by integrating forces over the moving body’s surface. Flow variables are second order accurate in space and first order accurate in time $(CFL<0.3)$. A least square cell-based gradient and SIMPLEC pressure-velocity coupling are used.
An explicit Volume of Fluid (VOF) approach in Equation 3 is adopted to capture the air-water interface using a Eulerian-Eulerian framework. In the continuity equation, the cells volume fraction $(0\leq\alpha_{q}\leq 1)$ of the fluid phase $q$ are included:
$\frac{\delta}{\delta t}\alpha_{q}+\triangledown(\alpha_{q}\bar{v_{q}})=\sum_{p=1}^{n}(\dot{m}_{pq}-\dot{m}_{qp})$ (3)
Surface tension effects are neglected, justified by a Weber number well in excess of unity and under-relaxation factors for momentum, viscosity, and body forces are reduced $(n<1)$ to stabilise convergence. Two additional equations are solved for capturing the turbulence in the flow by employing the $k-\omega$ SST turbulence model. A turbulence damping source term was added to limit the over prediction of viscosity near the water-air interface, which arises from the VOF formulation whereby the turbulence models tend to over predict eddy viscosity from the high strain rate at the interface [2].
Mesh and time-steps refinement studies were conducted to identify the coarsest discretisation that preserves solution accuracy.
Linear Hydrodynamic Reference Model
Decomposition of Hydrodynamic Forces
For a heaving body, the total vertical hydrodynamic force can be decomposed into hydrostatic $F_{hs}$, radiation $F_{rad}$ and excitation $F_{exc}$ contributions, the wave excitation contribution can further be decomposed into the Froude-Krylov and diffraction force components.
$F_{total}=F_{\mathrm{hs}}(x)+F_{\mathrm{rad}}(x,\dot{x})+F_{\mathrm{exc}}(t),$ (4)
where $x(t)$ denotes the heave displacement. Assuming linearity, this leads to the familiar forced damped-spring mass system in Equation 5
$(M+A)\ddot{x}(t)+C\dot{x}(t)+Kx(t)=F_{\mathrm{exc}}(t),$ (5)
with $M$ the structural mass, $A$ the added mass in heave, $C$ the radiation damping, and $K$ the hydrostatic stiffness i.e the buoyancy.
Linear Properties of a Heaving Sphere
For the spherical body used in this study, the structural mass and linear hydrostatic stiffness follow directly from the geometry,
$M=\frac{2}{3}\rho\pi r^{3},\qquad K_{linear}=\rho g\pi r^{2}$ (6)
where $\rho$ is the water density, $g$ is gravitational acceleration, and $r$ is the sphere radius.
The linear added mass $A$ and radiation damping $B$ coefficients for the heaving sphere are obtained analytically in the work of [13].
Heave decay Response of the Isolated Sphere
When the sphere is far from any wall and no incoming waves are present, the system reduces to a homogeneous spring–mass–damper oscillator driven only by linearly independent hydrostatics and radiation forces. The equation of motion becomes
$\bigl(M+A)\ddot{x}(t)+C\dot{x}(t)+Kx(t)=0$ (7)
The undamped natural frequency and (frequency-dependent) damping ratio are $\omega_{0}=\sqrt{\frac{K}{M+A}}$ and $\delta(\omega)=\frac{C}{2\bigl(M+A\bigr)}$ respectively. The resulting heave decay takes the standard damped-oscillator form $\sqrt{\omega_{0}^{2}-\delta(\omega_{d})^{2}}$ where $C_{1}$ and $C_{2}$ are set by the initial displacement and velocity.
This gives the general solution to the ODE
$x(t)=(C_{1}cos(\omega_{d})t+C_{2}sin(\omega_{d}t))e^{-\delta\omega_{d}t}$ (8)
For small drop heights, this linear model reproduces the CFD and experimental heave-decay envelopes with good accuracy (see Figure  4), confirming an adequate linear representation for small-amplitude motion. But, as the drop height increases, nonlinear hydrostatic and radiation effects become significant, and the linear decay solution progressively deviates from the reference CFD, with the linear model under-predicting the peak amplitude and the decay frequency.
Figure 4: Validation of drop test from drop height of (a) $0.1$ and (b) $0.5$ of the spheres diameter. Comparison between our CFD simulation, experimental data [14] and linear model.
Sphere in Regular Waves: Linear Excitation Model
When the sphere is subjected to regular incident waves, the motion can be captured in a linear ODE similar to that of a forced damped–spring–mass system. For a deep-water, first-order, linear Stokes wave theory, the free-surface elevation is given by:
$\eta(t)=A_{i}\cos(\omega t),$ (9)
and the corresponding wave-excitation force in heave is:
$F_{exc}(t)=A_{i}\cos(\omega t)X,$ (10)
where $A_{i}$ is the incident wave amplitude, and $\omega$ is the frequency of the incident waves. $X$ is the frequency dependent excitation force coefficient of a sphere, calculated using the commercial potential flow solver WAMIT in the work by [14].
The equation of motion becomes
$\bigl(M+A)\ddot{x}+B\dot{x}+Cx=F_{exc}(t)$ (11)
In a steady state, the linear solution has the form
$x(t)=x_{0}\cos(\omega t-\varphi),$ (12)
where $x_{0}$ is the response amplitude and $\varphi$ is the phase lag. The associated response amplitude operator (RAO) in heave is
$\mathrm{RAO}(\omega)=\frac{x_{0}}{A_{i}}$ (13)
Given analytically derived linear coefficients, the following linear model for the RAO is given in the paper by [18]:
$\mathrm{RAO}(\omega)=\frac{X}{\omega\sqrt{C^{2}+(\omega(M+A)-K/\omega)^{2}}}$ (14)
CFD and experimental RAOs are calculated in the validation study by taking the square root of the peak frequency identified in the power spectral density (PSD) of the heave response $S_{heave}(\omega)$ over that of the incident wave $S_{wave}(\omega)$.
$\mathrm{RAO}(\omega)=\sqrt{\frac{S_{heave}(\omega)}{S_{wave(\omega)}}}$ (15)
Figure 6 shows a comparison between the linear model and the CFD derived RAOs. It is evident that the CFD results systematically depart from the linear model (Equation 14) when the wave period is close to the spheres fundamental frequency and as wave steepness increase. This behaviour is consistent with the observations of [14], who showed that higher-order hydrodynamic effects become increasingly significant with larger wave steepness $(s)$. Correspondingly, the PSD in Figure 6 reveals a redistribution of energy from the fundamental frequency $\omega$ into higher harmonics (e.g. $2\omega$ and $3\omega$). As noted by [21], a linear time-invariant system cannot transfer energy between frequencies; thus, the appearance of these harmonics provides clear evidence of nonlinear dynamical behaviour.
Figure 5: Validation of RAO for our CFD by comparing our results for the spheres heaving RAO with a linear model
Figure 6: Normalised PSD of the sphere at resonance $T_{wave}=4.4$s for a steep wave $s=0.01$, a less steep wave $s=0.02$, and a linear response model
These discrepancies are indicative of nonlinear diffraction force effects. In the CFD simulations, the instantaneous wave field around the sphere is strongly modified by the large body motion, the free-surface elevation, local waterline geometry and scattered wave patterns. This, we hypothesise, generates higher-order components in the scattered wave system and alters the effective excitation force. As a result, the true excitation cannot be represented by a single linear force coefficient $X$ relating the free surface $\eta$ to the excitation force $F_{exc}$.
Non-linear Analytical Fluid Oscillator Model
Hydrodynamic Modelling for Launch and Recovery
For launch-and-recovery scenarios several effects undermine linear assumptions. Large relative motions, steep waves, and geometric complexity introduce; higher-order wave interactions, non-linear changes in hydrostatic restoring force due to the instantaneous submerged volume, and non-linear excitation forces. Previous CFD and experimental studies of heaving bodies in waves have shown that these mechanisms lead to significant deviations from linear theory [18, 24], wherby excitation and radiation forces are no longer separable by simple superposition.
These observations motivate the present reduced-order modelling strategy, where, rather than abandoning the low-dimensional ODE representation, we treat its coefficients as nonlinear functions of both the system-state and the input-space $(WD,DH)$, and infer those functions directly from CFD data.
A starting point is therefore to represent the dynamics using the following non-linear ODE of the form
$(M+A)\,\ddot{x}+C(x,\dot{x})\,\dot{x}+K(x)\,x=F_{\mathrm{exc}}(t),$ (16)
where the effective added mass $A$, damping $C$, stiffness $K$, and excitation $F_{\mathrm{exc}}$ can be derived either analytically from first principles or from the raw data using linear regression based techniques.
Nonlinear Hydrostatic Restoring Force
For the capped sphere in Figure 7, the exact hydrostatic restoring force can be written in terms of the submerged cap height $h$ such that;
$V(h)=\frac{\pi}{3}h^{2}(3r-h),\qquad h=DH\,x+r,$ (17)
where $R$ is the sphere radius, $DH$ is the initial drop height, and $x$ is the non-dimensional heave displacement (scaled by $DH$).
Figure 7: Capped sphere geometry
The vertical force is then normalised by drop height and mass in:
$\hat{F}=\frac{F}{(M+A)\,DH}=\frac{V\rho g}{(M+A)DH},$ (18)
which, after substitution of h=DHx+rh=DH\,x+r and expansion, yields a hydrostatic relationship with a linear and cubic dependence on $x$. Writing the restoring term in the generic polynomial form
$k(x,x^{2},x^{3})x=\Xi_{1}x+\Xi_{3}x^{2}+\Xi_{5}x^{3},$ (19)
and matching coefficients gives the analytical approximation
$k(x,x^{2},x^{3})\,x\approx 2.066\,x-0.02755\,DH^{2}x^{3}.$ (20)
The corresponding non-dimensional hydrostatic restoring curve is shown in Figure 8, illustrates the departure from the linear approximation.
Figure 8: Non-dimensional hydrostatic restoring force for the capped sphere, showing the cubic hardening behaviour with increasing displacement.
Numerical Time Integration of Nonlinear ODEs
To assess the predictive capability of our nonlinear analytical model, the ODE is integrated in time and the resulting heave–decay response is compared against the CFD trajectories. The second–order equation
$\ddot{x}(t)=f\bigl(x,\dot{x},t\bigr)$ (21)
is recast as a first–order system by introducing the state vector
$\mathbf{s}(t)=\begin{bmatrix}x(t)\\[2.0pt] \dot{x}(t)\end{bmatrix},\qquad\dot{\mathbf{s}}(t)=\mathbf{f}\bigl(\mathbf{s}(t),t\bigr)=\begin{bmatrix}\dot{x}(t)\\[2.0pt] f\bigl(x(t),\dot{x}(t),t\bigr)\end{bmatrix}$. (22)
One explicit fourth–order Runge–Kutta (RK4) step from $t_{n}$ to $t_{n+1}=t_{n}+\Delta t_{n}$ is then given by
$\mathbf{k}_{1}$ $=\mathbf{f}\bigl(\mathbf{s}_{n},t_{n}\bigr),$ (23)
$\mathbf{k}_{2}$ $=\mathbf{f}\bigl(\mathbf{s}_{n}+\tfrac{\Delta t_{n}}{2}\mathbf{k}_{1},t_{n}+\tfrac{\Delta t_{n}}{2}\bigr),$ (24)
$\mathbf{k}_{3}$ $=\mathbf{f}\bigl(\mathbf{s}_{n}+\tfrac{\Delta t_{n}}{2}\mathbf{k}_{2},t_{n}+\tfrac{\Delta t_{n}}{2}\bigr),$ (25)
$\mathbf{k}_{4}$ $=\mathbf{f}\bigl(\mathbf{s}_{n}+\Delta t_{n}\,\mathbf{k}_{3},t_{n}+\Delta t_{n}\bigr),$ (26)
$\mathbf{s}_{n+1}$ $=\mathbf{s}_{n}+\frac{\Delta t_{n}}{6}\left(\mathbf{k}_{1}+2\mathbf{k}_{2}+2\mathbf{k}_{3}+\mathbf{k}_{4}\right),$ (27)
In all simulations a uniform time step of $\Delta t=0.02\,\mathrm{s}$ is used, corresponding to $N=1500$ steps, or $30\,\mathrm{s}$ in real time over the decay window.
Comparison with Regression-based Nonlinear Model
Unlike the linear case, for which a closed-form singular, smooth analytical solution exists, the inclusion of nonlinear hydrostatic and damping terms generally precludes a simple explicit solution to the ODE. Instead, the analytically derived nonlinear restoring and damping terms are compared with those obtained by regression (SINDy-style least squares) from the CFD time series.
For a representative large drop height of $DH=5\,\mathrm{m}$, the analytical model for the dominant terms in the acceleration reads
$\ddot{x}_{\mathrm{analytical}}(DH=5\,\mathrm{m})=-2.063\,x+0.689\,x^{3}-0.24\,\dot{x},$ (28)
whilst the corresponding regression-based model obtained from the raw CFD signal is
$\ddot{x}_{\mathrm{regression}}(DH=5\,\mathrm{m})=-2.057\,x+0.686\,x^{3}-0.148\,\dot{x}.$ (29)
The close agreement in the $x$ and $x^{3}$ coefficients confirms that the nonlinear hydrostatic contribution is well captured analytically.
Comparison between Non-linear Analytical and Numerical
Figure  9 compares the numerically integrated nonlinear model with the raw CFD data for two representative drop heights. For small drops ($DH=1\,\mathrm{m}$), the response remains close to linear and the nonlinear corrections are modest, whereas for larger drops ($DH=5\,\mathrm{m}$) (Equations 28 and 29) the cubic hydrostatic term and amplitude-dependent damping play a dominant role in shaping the decay envelope and improving agreement with the CFD solution.
Figure 9: Heave decay for (a) $DH=5\,\mathrm{m}$ and (b) $DH=1\,\mathrm{m}$ comparison between raw CFD data and the nonlinear analytical ODE model integrated with RK4.
Sparse Identification of Nonlinear Dynamics (SINDy)
Let’s postulate that the normalised heave response satisfies a single-degree-of-freedom nonlinear fluid oscillator of the form
$\ddot{x}(t)=\sum_{k=1}^{6}\Xi_{k}\,\ddot{x}_{k}\!\left(x,\dot{x}\right)+\Xi_{7}\sin(\omega_{0}t)+\Xi_{8}\cos(\omega_{0}t),$ (30)
SINDy is the framework by which an ODE model can be recovered directly from time series data. $\ddot{x}$ can be represented as a sparse linear combination of candidate non-linear functions, allowing for the assembly of a candidate library matrix for the linear and non-linear terms forming our ODE;
$\Theta=\begin{bmatrix}x&\dot{x}&x^{2}&\dot{x}^{2}&x^{3}&\dot{x}^{3}&\sin(\omega t)&\cos(\omega t)\end{bmatrix}$
and seek a coefficient vector $\Xi$ such that
${\Xi}={\arg\min}\;\bigl\|\mathbf{x}-{\Phi}{\Xi}\bigr\|_{2}^{2}+\lambda\bigl\|{\Xi}\bigr\|_{2}^{2},$ (31)
where $\lambda$ is a regularisation parameter. This system is solved by forming the equations in ridge regression such that
$\bigl({\Phi}^{T}{\Phi}+\lambda I\bigr){\Xi}={\Phi}^{T}\ddot{x}$ (32)
a pseudoinverse $({\Phi}^{T}{\Phi}+\lambda I\bigr)^{-1}$ is taken based on singular value decomposition (SVD) $\Phi=U\Sigma V^{T}$ to solve for the coefficients $\Xi$.
Sparsity is then enforced via Sequential Thresholded Least Squares (STLSQ) [12]. At each iteration $k$:
  1. 1.
    Identify “small” coefficients for $\Xi_{k}<\lambda$ where $\lambda$ is a sparsity threshold.
  2. 2.
    Set ${\Xi}_{k}=0$ and refit Equation 31 using only the remaining “active” columns of $\Phi$}.
  3. 3.
    This process is repeated until the active set stabilises.
The coefficients are subsequently mapped back to the original, unnormalised feature space, yielding a sparse vector $\Xi$ associated with the functions in Equation 30.
Here, the SINDy-identified terms can be grouped into physically meaningful contributions. For example, the polynomial terms represent nonlinear hydrostatic and radiation forces, while the sinusoidal entries are a crude representation of the wave excitation force arising from the interaction between the sphere and the rebounded waves.
To determine an appropriate candidate library, SINDy is applied separately to each $(WD,DH)$ case by solving a regularised least-squares problem and evaluated by taking the average MSE over the entire dataset. Figure 10 shows the candidate library evaluated in a for-loop over increasing polynomial and harmonic orders, which allows us to identify the best set of terms required to obtain a satisfactory representation of the dynamics
Figure 10: The required non-linear terms include in the library to best capture the dynamics (MSE).
After evaluating the accuracy of several candidate libraries, the optimal model was identified as the third-order polynomial plus first-order harmonic library in Equation 30. This basis is both compact and sufficiently accurate, achieving an error of $MSE=5.1\times 10^{-4}$. Including higher-order harmonics did not yield any improvement in accuracy; a better model would require a more realistic the excitation-force model.
Excitation Force Model
To represent the wave excitation forces from the waves rebounding from the wall, the SINDy feature library is augmented with single–frequency harmonic terms in Figure 11, such that:
$\Theta\;\supset\;\begin{bmatrix}\sin(\omega t)&\cos(\omega t)\end{bmatrix}$,
which act as a simple, low–order surrogate for the excitation force.
Figure 11: Simplified excitation force model
From the identified coefficients of $\sin(\omega t)$ and $\cos(\omega t)$, SINDy provides an estimate of the effective excitation amplitude and phase. This estimate is useful in capturing the general magnitude and phase of the excitation force, but as it remains a crude approximation, it is incapable of fully capturing the true physics.
In an attempt to resolve this issue, let’s begin by assuming that excitation is linearly separable from hydrostatic and radiation effects, so that the normalised excitation force $\hat{F}=\frac{F}{(M+A)DH}$ can be obtained as
${\hat{F}}_{\mathrm{exc}}={\hat{F}}_{\mathrm{total}}-{\hat{F}}_{\mathrm{homogeneous}},$
leading to the reconstructed excitation shown in Figure 12. Under this assumption, the homogeneous dynamics for a representative case (e.g. $DH=5\,\mathrm{m}$) can be written as
$\ddot{x}+2.057\,x-0.686\,x^{3}+0.148\,\dot{x}=0,$ (33)
and one might hope to write a non-linear ODE representing the dynamics of the sphere with the returning waves from the wall as
$\ddot{x}+2.057\,x-0.686\,x^{3}+0.148\,\dot{x}=F_{\mathrm{exc}}(t).$ (34)
such that
$F_{\mathrm{exc}}(t)=X\eta(t)$ (35)
Figure 12: Linear excitation forces assumption
In practice, this linear superposition fails such that:
$(M+A)\ddot{x}+C(x,\dot{x})\dot{x}+Kx\neq X\eta(t)$ (36)
The excitation force $F_{\mathrm{exc}}(t)$ inferred in this way does not reproduce the CFD heave time history when re–integrated, and the true excitation force $F_{exc}$ required to recover the true trajectory is shown in Figure 13 differs markedly from Figure 12. This indicates that excitation force is non-linearly coupled to the body motion and radiation field, so that the true excitation cannot be represented by a single linear term that is independent of the state. In future work, explicitly considering the Froude-Krylov and radiation forces separately may be considered to potentially provide an improved model of the excitation forces.
Figure 13: True excitation forces
Despite this lack of linearity, there is potentially some information on the excitation force that can be recovered in the simplified $sin/cos$ coefficients determined using our SINDy regression.
The phase of the excitation force can be predicted from a simple first-principles derived wave travel time argument based on the WD and the wave celerity $(c)$:
$t_{\mathrm{delay}}=\frac{2WD}{c},\quad\phi=\omega t_{\mathrm{delay}},$ (37)
and matches well with the phase extracted via SINDy in Figure 14.
Figure 14: SINDy identified excitation force phase vs analytical predictions
SINDy identifies sinusoidal excitation terms whose phase aligns closely with the wave-travel-time prediction. Around WD=15 m, a minimal excitation forces amplitude is observed (Figure 15), which is the same distance where the wave is 90 degrees offset from the spheres motion. No simple analytical relation for the excitation force magnitude has been found, but would likely require a similar methodology as [15].
Figure 15: SINDy identified excitation force magnitude
SINDy-informed global neural operator
Sparse regression provides an interpretable ODE for each individual $(WD,DH)$ case, but the corresponding coefficient vectors $\Xi$} are only available at discrete points in the two-dimensional parameter space. The aim is to achieve a global surrogate valid across the full $(WD,DH)$ domain by approximating the mapping
${\Xi}={\Xi}(WD,DH)$ (38)
with a neural operator network (ONet). The key design choice is that the network is trained to match the trajectory $x(t)$ rather than the coefficients themselves (Equation 42), reflecting the non-uniqueness of nonlinear ODE representations, as multiple different coefficient matrices ($\Xi$}) can generate almost identical dynamics.
Figure 16 shows a 2D slice for the magnitude of the SINDy-derived coefficients across all wall distances. This information is used as a prior within the ONet to restrict the results to physically plausible ranges for each component of $\Xi$}.
Figure 16: SINDy-identified coefficients.
The ONet is implemented as a fully connected feed-forward network such that;
$\mathcal{N}_{\theta}:\mathbb{R}^{2}\xrightarrow{NN(WD,DH)}\mathbb{R}^{8}$ (39)
The network architecture used;
$2(WD,DH)\xrightarrow{}32(W\times x^{l-1}+b_{i})\xrightarrow{tanh(\cdot)}8(\Xi_{normalised})$ (40)
consists of a single hidden layers of width 32 and a $\tanh$ non-linear activation. The final $\tanh$ layer yields output values from the neural networks such such that;
${{\Xi}_{normalised}}\in(-1,1)^{8},$
Figure 17: SINDy-identified coefficients normalised between (1,1)(-1,1).
After the output from the NN, in Figure 17, the values are rescaled using SINDy-derived bounds:
$\Xi_{i}=\frac{\xi_{\max}^{i}-\xi_{\min}^{i}}{2}\,{\Xi}_{i}+\frac{\xi_{\max}^{i}+\xi_{\min}^{i}}{2},\qquad i=1,\dots,8.$ (41)
Here $(\xi_{\min}^{i},\xi_{\max}^{i})$ are chosen so that the resulting $\Xi_{k}$ span the range of SINDy-identified coefficients over all CFD trajectories. Equation (41) therefore acts as a prior, constraining the learned coefficients to plausible hydrodynamic values.
Given a set of $N_{\mathrm{tr}}$ CFD trajectories, the ONet parameters $W_{k},b_{k}$ are obtained by minimising the mean-squared error between the predicted and CFD responses,
$MSE_{loss}=\frac{1}{N_{\mathrm{tr}}}\sum_{i=1}^{N_{\mathrm{tr}}}\frac{1}{T_{\mathrm{s}}}\sum_{j=0}^{T_{\mathrm{s}}-1}\left[x_{\theta}^{(i)}(t_{j})-x^{(i)}(t_{j})\right]^{2}$ (42)
where $x_{\theta}^{(i)}(t_{j})$ is obtained by numerically integrating using RK4 (Equation 27) the nonlinear ODE with coefficients ${\Xi}=\mathcal{N}_{\theta}(WD^{(i)},DH^{(i)})$. The loss (42) is minimised using the Adam optimiser.
The test loss is monitored to detect overfitting. Once trained, the ONet can be evaluated at arbitrary $WD,DH$ values, providing an efficient surrogate for the CFD solver.
In summary, this framework involves local sparse regression based ROMs followed by global operator network based interpolation. SINDy enforces interpretability, while the neural operator supplies a smooth, data-driven coefficient manifold over parameter space (Figure 20), yielding a real-time physics-constrained surrogate model for the heave decay near a wall.
Training
The network training utilises a standard PyTorch optimisation loop. However, the computational efficiency is constrained by the nature of the loss function, which requires numerical integration of the governing differential equations. The numerical integration is implemented sequentially, preventing effective parallelisation on a GPU. Consequently, the forward and backward passes are executed primarily in series on the CPU, resulting in a training duration of approximately 24 hours.
Careful parameter initialisation was considered, as an essential first step to training the algorithm to ensure stability. In this, we stochastically initialise weights and biases but restrict the range of possible coefficient values when de-normalising. These constraints are gradually relaxed as training progresses, ensuring numerical stability over the fist few training epochs.
Figure 18: Training data (orange) and testing data (green).
Results
Surrogate Accuracy
The maximum achievable accuracy over the domain, based on the idealised SINDy-derived coefficients, is approximately $5.1\times 10^{-4}$ (see Figure 10). In comparison, the neural operator achieves an average training and testing (interpolation-set) error of
$\mathrm{MSE}_{\mathrm{train}}\approx 4.85\times 10^{-4},\qquad\mathrm{MSE}_{\mathrm{test}}\approx 5.76\times 10^{-4}.$
Here, the reported testing error excludes the extrapolated case at $WD=25.0\,\mathrm{m},\ DH=1.30\,\mathrm{m}$, which lies outside the training domain in Figure 18.
Interestingly, both the training and interpolation errors fall slightly below the theoretical SINDy lower bound. In conjuction with the fact that the NN identified manifolds do not itersect the SINDy identified coefficients, this indicates that the neural operator is not replicating the linear-algebra-derived coefficients, but is instead learning a different representation that yields trajectories more consistent with the raw data.
Figure 19: Neural-operator prediction vs CFD for test cases.
Figure 20: Learned coefficient manifolds
Discussion
Interpolation and Extrapolation
Figure 18 shows which validation and test points lie within the interpolation region and which correspond to extrapolation. As expected, interpolation yields more accurate trajectories, while extrapolated conditions produce noticeably larger errors.
NN Outputs
The neural network–identified coefficient manifolds are shown in Figure  20. These surfaces represent the NN’s prediction throughout the entire input space with the training dataset, represented by the blue markers, lying on this manifold. This confirms that the NN has successfully learned a smooth parametric representation of the coefficient.
The neural operator’s smoothed coefficient manifolds deviates significantly from those obtained by direct linear algebra through SINDy in Figure 20. However, both sets generate nearly identical dynamical responses. This brings a key insight whereby we hypothesise that this is due to the fact that for nonlinear systems, there is no singular best solution. Note that the neural operator identifies a set of coefficients that minimises trajectory error rather than the error to the SINDy derived coefficient.
Conclusion
The results demonstrate that interpretable nonlinear ODEs can be reconstructed from CFD data using sparse regression, and that these local models can be embedded into a global operator network to form an efficient reduced-order surrogate. Across the sampled input domain, the SINDy-based identification consistently recovers hydrostatic and radiation contributions that agree with analytical expectations and with previously validated linear and weakly nonlinear models for the heaving sphere. In particular, the close agreement between analytically derived and regression-based coefficients for the linear and cubic stiffness terms confirms that the dominant nonlinear hydrostatic mechanisms are well captured. This gives confidence that the learned surrogate is anchored to physically meaningful force components rather than arbitrary empirical fits.
By contrast, the excitation terms associated with reflected waves are found to be substantially more complex. The failure of a linear decomposition strategy to isolate excitation force may indicate that the excitation force may not be treated as state-independent forcing. The simple single-frequency $\sin/\cos$ terms included in the SINDy library provide only a crude surrogate for these effects. They recover the correct phase delay, which is consistent with a wave travel-time argument based on $2WD/c$, but they do not reproduce the true excitation force realistically. This underscores the need for more sophisticated forcing representations if diffraction-dominated regimes are to be modeled with higher fidelity, and allow the learned coefficients to reproduce a higher accuracy representation of the data.
The SINDy-informed neural operator offers an interpretable method of designing the surrogate model. Rather than attempting to learn the dynamics directly from trajectories in a black-box fashion, the operator acts on a low-dimensional space of coefficients.
The analysis on interpolation and extrapolation further illustrate the strengths and limitations of the approach. Within the training data, the network attains mean-squared errors close to the theoretical lower bound imposed by the SINDy-based ideal coefficients, indicating that the mapping $(WD,DH)\mapsto{\Xi}$ has been learned with high accuracy.
From an applications perspective, the proposed ROM provides a useful building block for launch-and-recovery analysis. It replaces expensive CFD runs with a real-time surrogate that can evaluate new $WD,DH$ configurations, while retaining a clear link between the model coefficients and underlying hydrodynamic mechanisms (added mass, damping, stiffness, and excitation). Moreover, the trajectory-based loss and physics-informed rescaling of coefficients act as priors that improve robustness relative to purely data-driven recurrent models.
At the same time, several open questions remain. The present work considers only a highly idealised geometry and a single mode of motion (heave), and it focuses on regularised decay responses rather than fully irregular seas or six-degree-of-freedom dynamics. Extending the methodology to more complex hull forms, coupling modes of motion will require both richer libraries of candidate nonlinear terms and more robust operator architectures. Additionally, while the current formulation uses CFD as the sole source of training data, the framework is compatible with experimental measurements, which could be incorporated in future.